OpenStudy (mtalhahassan2):

Need help!! Diameter of a sphere is measure in 2.14 micro- meter. Calculate the circumference and volume of a sphere in the meter.

OpenStudy (anonymous):

Convert 2.14 micrometer to meters. Find the radius by dividing the diameter by 2 Use it for the formula for the volume and circumference of a sphere. Circumference of a sphere=2(pi)r Volume of a sphere=\[\frac{ 4 }{ 3 }pir^3\]

OpenStudy (mtalhahassan2):

Circumference of a sphere=2(pi)r =2 (3.14)(2.14) =1.34

OpenStudy (anonymous):

Like I said: 1. Convert 2.14 micrometer to meters. 2. Find the radius by dividing the diameter by 2

OpenStudy (anonymous):

\[Diameter=\frac{ radius }{ 2 }\]  1 meter=1.0 x 10^6 micrometers so 1 micrometer=1.0 x 10^-6 meters

OpenStudy (mtalhahassan2):

then the circumference be =2(3.14)(1.07) =6.7

OpenStudy (mtalhahassan2):

@Shalante

OpenStudy (mtalhahassan2):

v= 4/3 pie r^3 v= 4/3(3.14)(1.07)^3

OpenStudy (anonymous):

Convert!!

OpenStudy (anonymous):

Convert to micrometer

OpenStudy (anonymous):

1 meter=1.0 x 10^6 micrometers so 1 micrometer=1.0 x 10^-6 meters

OpenStudy (mtalhahassan2):

oh ok then it be 6.72 x 10^-6 meters

OpenStudy (mtalhahassan2):

rightz?

OpenStudy (mtalhahassan2):

but know for the volume

OpenStudy (mtalhahassan2):

is it be 4*pi*1.07^3

OpenStudy (mtalhahassan2):

sorry it be 4/3*pi*1.07^3

OpenStudy (anonymous):

Hey, my explanation above says Convert 2.14 micrometer to meters. Find the radius by dividing the diameter by 2 Use it for the formula for the volume and circumference of a sphere. Circumference of a sphere=2(pi)r Volume of a sphere=(4/3)(pi)r^3 You convert first before plugging anything or you will get messed up because there is an exponent.

OpenStudy (anonymous):

Do the steps in order please.

OpenStudy (anonymous):

\[\frac{ micrometers }{ 1 }\times \frac{ 0.00001 }{ 1micrometer }\] Convert to meters* from above.

OpenStudy (mtalhahassan2):

ok I get the answer

OpenStudy (anonymous):

Show me your steps and answer.

OpenStudy (mtalhahassan2):

ok wait a sec

OpenStudy (mtalhahassan2):

V=43πr3 V=43π(1.07×10−6 m)3 V=43π 1.073×10−6⋅3 m3 V=5.13×10−18 m3

OpenStudy (anonymous):

(4/3)(pi)^3 you mean?

OpenStudy (mtalhahassan2):

yeah

OpenStudy (anonymous):

I got 5.12 x 10^-18 but close enough.

OpenStudy (anonymous):

And your first one?

OpenStudy (mtalhahassan2):

that is C= 2*pi*1.07 um = 6.7x 10^-6

OpenStudy (anonymous):

Yea, it looks correct

OpenStudy (mtalhahassan2):

ok thanks a lot